OpenStudy (anonymous):

evalu(ate the integral sqrt(4+x^2)/(X^4)dx from 0 to 1

OpenStudy (anonymous):

if we rationalize the numerator .. maybe that will get it into a more familiar form

OpenStudy (anonymous):

nah, that dint work out nicely

OpenStudy (anonymous):

let u = 4+x^2, x^2 = u-4; 2x dx = 1 du sqrt(u) du ------ -------- (u-4)^2 2sqrt(u-4) hmm

OpenStudy (anonymous):

exact or approximate?